OpenStudy (1645323):

Find the focus of the parabola y = ½(x - 4)2 - 3 Select one: a. (3.5, -3) b. (4, -3.125) c. (4, -2..875) d. (4.5, -3)

OpenStudy (1645323):

@terenzreignz

OpenStudy (tkhunny):

x-coordinate is 4 -- obvious by inspection (axis of symmetry). y-coordinate > -3 -- obvious by inspection (positive leading coefficient with vertex at (4,-3))

OpenStudy (1645323):

@terenzreignz

OpenStudy (1645323):

@TuringTest

OpenStudy (1645323):

@Luis_Rivera @zepdrix

OpenStudy (1645323):

How do you solve this @jdoe0001

OpenStudy (jdoe0001):

\(\bf 4p(y-{\color{blue}{ k}})=(x-{\color{red}{ h}})^2\qquad \begin{array}{llll} vertex\ ({\color{red}{ h}}\quad ,\quad {\color{blue}{k}})\\ p=\textit{distance from the vertex to the focus} \end{array} \\------------------------------- \\ \quad \\ y=\cfrac{1}{2}(x-4)^2-3\implies y+3=\cfrac{1}{2}(x-4)^2\implies (y+3)=\cfrac{(x-4)^2}{2} \\ \quad \\ 2(y+{\color{blue}{ 3}})=(x-{\color{red}{ 4}})^2\) so.. what do you think "p" is?

OpenStudy (tkhunny):

What, thinking about the definition was too easy?

OpenStudy (jdoe0001):

eheh

OpenStudy (1645323):

what is p?

OpenStudy (jdoe0001):

well, what is it? notice the template and what you'd have \(\bf {\color{green}{ 2}}(y+{\color{blue}{ 3}})=(x-{\color{red}{ 4}})^2\\ {\color{green}{ 4p}}(y-{\color{blue}{ k}})=(x-{\color{red}{ h}})^2\) what do you think would be "p"?

OpenStudy (1645323):

1

OpenStudy (jdoe0001):

4p = 2

OpenStudy (1645323):

p=2

OpenStudy (jdoe0001):

so find "p" and from the vertex, move over THAT MUCH to get the point of the focus

OpenStudy (1645323):

ok so 4,-1?

OpenStudy (1645323):

the vertex is 4,-3

OpenStudy (jdoe0001):

hmmm what did you get for the vertex?

OpenStudy (jdoe0001):

right so .... what's "p" ?

OpenStudy (1645323):

p is 2

OpenStudy (jdoe0001):

4p = 2 so if you solve for "p" you get 2 ?

OpenStudy (1645323):

no nevermind .5

OpenStudy (1645323):

so the foucs is 3.5,-3

OpenStudy (jdoe0001):



OpenStudy (1645323):

ok thanks a lot

OpenStudy (jdoe0001):

yw

OpenStudy (tkhunny):

?? y = ½(x - 4)2 - 3 Vertex: (4,-3) Axis of Symmetry: x = 4 x is quadratic and y is linear (Opens up or down) 1/2 > 0 (Opens up) Focus: (4,-3+Something) Directrix: y = -3-Something How did the focus move from x = 4?